OpenStudy (anonymous):

Which ordered pair is a solution of the inequality? y4x - 5 A. (3, 4) B. (2, 1) C. (3, 0) D. (1, 1)

OpenStudy (imstuck):

What sign is the inequality sign and where does it go in your expression?

OpenStudy (aum):

y4x - 5 does not look right. Make correction if necessary.

OpenStudy (anonymous):

nvm What is the solution of the system? Use substitution. y = x + 6 y = 2x A. (6, 12) B. (–12, –6) C. (–6, –12) D. (2, 4)

OpenStudy (aum):

y = x + 6 y = 2x Put y = 2x in the first equation and solve for x. Then solve for y.